Hahn Company uses a job-order costing system. Its plantwide predetermined overhead rate uses direct labor-hours as the allocaplease also calculate the plantwide predetermined overhead rate
Answer #1

1) Direct labor hour used in Job Omega = 401450/15.50 = 25900

Predetermine overhead rate = 220150/25900 = 8.50 per hour

2) Calculate following

Job Alpha
Direct material 244400
Direct labor (64400*15.50) 998200
Manufacturing overhead applied (64400*8.50) 547400
Total Job cost 1790000